OpenStudy (anonymous):

How to show T is a linear operator from P^2 to P^2

OpenStudy (anonymous):

depends largely on \(T\)

OpenStudy (anonymous):

Yea forgot (T_p)(x) = p(x+1)

OpenStudy (anonymous):

what is \(P^2\) ?

OpenStudy (anonymous):

Space of polynomials having no more than degree 2 and coefficients in R

OpenStudy (anonymous):

Basis of P^2 is (1,x,x^2)

OpenStudy (anonymous):

Just not sure how to show T(f+g) = Tf +Tg for

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

polynomial of degree two looks like \[a_2x^2+a_1x+a_0\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

take two general ones, say \[a_2x^2+a_1x+a_0\] and \[b_2x^2+b_1x+b_0\] add them, get \[(a_2+b_2)x^2+(a_1+b_1)x+(a_0+b_0)\] then take \[T((a_2+b_2)x^2+(a_1+b_1)x+(a_0+b_0))\]

OpenStudy (anonymous):

once you do that, hold on to that though, then compute \[T(a_2x^2+a_1x+a_0)+T(b_2x^2+b_1x+b_0)\] and see if you get the same thing

OpenStudy (anonymous):

Yea cool. Having trouble getting what the right side will look like using (T_p)(x) = p(x+1). What you said looks right but I think you have to use the T operator I mentioned

OpenStudy (anonymous):

This should be enough though

OpenStudy (anonymous):

i am a bit confused as to what \(p(x+1)\) actually means

OpenStudy (anonymous):

it is the polynomial at \(x+1\) ?

OpenStudy (anonymous):

It is p(x+1) where p is a polynomial but it is not (p)(x+1), ((p times (x+1)

OpenStudy (anonymous):

if P = standard 2nd degree polynomial not sure how to write it in using p(x+1)

OpenStudy (anonymous):

if so \[T(a_2x^2+a_1x+a_0)=a_2(x+1)^2+a_1(x+1)+a_0\]

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

I was stuck or something but that is right, thanks a bunch!! saved me